The angular velocity of an airplane propeller increases from 12.0 rad/s to 16.0 rad/s while turning through 7.00 rad. What is the angular acceleration in rad/s^2?


1
Expert's answer
2021-11-16T09:53:44-0500

The angular acceleration is


"\alpha=\frac{\omega_1^2-\omega_2^2}{2\phi}=8\text{ rad/s}^2."